OpenStudy (anonymous):

Bobby is testing the effectiveness of a new cough medication. There are one hundred people with a cough in the study. Seventy patients received the cough medication and thirty other patients did not receive treatment. Thirty-four of the patients who received the medication reported no cough at the end of the study. Twenty of the patients who did not receive medication reported no cough at the end of the study. What is the probability that a patient chosen at random from this study took the medication, given that they reported no cough?

OpenStudy (anonymous):

I keep getting 69.4% which is not an answer choice

OpenStudy (anonymous):

I did P(A)=.7 P(A&B)=.486 P(B|A)=.486/.70=.694

OpenStudy (chris911):

refer to this one to help u http://openstudy.com/study#/updates/5277f8aee4b0ebf890293ca8

OpenStudy (chris911):

scroll all the way down

OpenStudy (phi):

Thirty-four of the patients who received the medication reported no cough Twenty of the patients who did not receive medication reported no cough so 34+20 = 54 reported no cough the probability of receiving the med is 34 out of 54

OpenStudy (phi):

using formulas, label A as taking med, B as "no cough" \[ P(A | B) = \frac{P(A \cap B)}{P(B)}\] Pr(B) = Pr(no cough) = 54/100= 0.54 Pr of reporting no cough and taking the med is 34/100 = 0.34

OpenStudy (anonymous):

Thanks @chris911 and @phi I got 62.96% as my final answer. Is that correct?

OpenStudy (phi):

yes, that is what I get.

OpenStudy (anonymous):

Thank you :)

OpenStudy (phi):

yw

OpenStudy (chris911):

yep